OpenStudy (yumeybitaa):

Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. -9 -3 3 9

OpenStudy (neer2890):

put x=3 and y=-2 in x-3y so x-3y=3-3(-2)=3+6=9